I am having a problem with this question. This guy took a trip two times. The first time he was charged $210 for 625 miles. The next time he was charged S133.5 for 370 miles. Whats is the base rate, and how much he was charged for each mile? What is the slope?
Answer by mananth(16949)

let the base rate be x
the mile rate be y/mile
..
x+625y=210
x+370y= 133.5
..
subtract the two equations
x+625y-x-370y=210-133.5
255y=76.5
y=76.5/255
=$ 0.3/mile
..
plug the value of y in any eauation
x+625*0.3=210
x=210-187.5
x=22.5 base rate
write the equation in the form of y=mx+c and compare the value with m